OpenStudy (anonymous):

if we have that x+2y=40 then the largest possible value for xy is what? maximum product = ?

OpenStudy (anonymous):

you remeber that fence problem? this can be done in the similar fashion.

OpenStudy (anonymous):

differenciate x+27=40 wrt x. You have 1+2dy/dx=0 that is dy/dx=-1/2 Now the max. product P=xy is constant so, differenciating that 0=x(dy/dx)+y that is, x(-1/2)+y=0 that is 2y=x

OpenStudy (anonymous):

btw, it's not x+27, it's x+2y (pardon that typo)

OpenStudy (anonymous):

Thus, y is 40/3 and x is 80/3 the product is 3200/9